OpenStudy (anonymous):

Can someone please help me with -1/4(x-12)+1/2(x+2)=x+4 , Linear Equations!

OpenStudy (lilg132):

ok i like to use numbers instead of fractions so first of all -1/4 = -0.25 1/2 = 0.5 so -1/4(x-12)+1/2(x+2)=x+4 would be -0.25(x - 12) + 0.5(x + 2) = x + 4 which is -0.25x + 3 + 0.5x + 1 = x + 4 put it in order 0.5x - 0.25x + 1 + 3 = x + 4 which is 0.25x + 4 = x + 4 now we rearrange to bring the numbers to one side and the variables to the other side which gives you 0.25x - x = 4 - 4 0.25x - x = 0 -0.75x = 0 anything multiplied by 0 will always give the answer 0 hence why x is 0 x = 0 it cant be any other answer dont forget to click good answer :)